# DTG : Diffusion-based Trajectory Generation for Mapless Global Navigation

### 14 Mar 2024 v1

@@ III-C Training Strategy
According to the problem defined in Section III-A, we have two targets to achieve: 1. reduce the distance between the generated trajectories and the ground truths and 2. minimize the portion of generated trajectories in non-traversable areas. Since our approach is an end-to-end model, we jointly train both targets with an adaptive training strategy:
@@
where $t\in\left\{1,...,N\right\}$ and $\epsilon_{d}\in\mathcal{N}(0,\mathbf{I})$. $\bm{\tau}_{gt}$ is the ground truth trajectory, because optimizing VDM [vdm] boils down to predict the original ground truth. SNR$(t)=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}$. $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ is defined the same as [ho2020denoising]. To simplify the loss function, because SNR(t) is monotonically decreasing, $(\text{SNR}(t-1)-\text{SNR}(t))$ is always positive. To simplify the training, we find similar training results after removing this term. Thus, the loss function is simplified as:
@@ III-C2 Adaptive Training of Diffusion Models
where $d(\cdot)$ calculates the distance between the current waypoint $w_{m}$ in $\hat{\bm{\tau}}_{t}$ and its nearest non-traversable area The distance function clips the value to $[0-1]$ meters. We sample t from the diffusion step buffer $\mathbf{b}_{t}$, which stores the currently available diffusion steps. As shown in the algorithm 1, $i$ is the index of training epochs and $N_{t}=10$ is the maximum step number to apply the traversability loss. $l_{d}(t-1)$ calculate the average of $5$ last recorded diffusion loss values at step $t-1$. If the average loss value is smaller than the threshold $H_{d}$, we add this step to the step buffer $\mathbf{b}_{t}$. We increase the range of diffusion steps incrementally to adapt the training of traversability to the training of ground truth trajectories. Finally, we have the total loss function for DTG: $\mathcal{L}=\mathcal{L}_{d}+\beta\mathcal{L}_{t}$, where $\beta$ is a hyperparameter.
@@
Figure 3: Traversability Analysis in Challenging Occluded Environment: The top row shows the generated trajectories (red) in camera view. The bottom row shows the top-down view in the traversability map. The cyan color represents the generated trajectories and the yellow color represents the most heuristic trajectory to the goal. DTG can generate trajectories w.r.t. the geometric shape of the traversable areas, but other approaches cannot generate trajectories fully traversable, where the non-traversable parts are marked by red circles.
@@
Figure 4: Travel-Distance Analysis in Challenging Narrow Passage Environment: The goal is behind the build. DTG can generates trajectory with higher distance ratio in a narrower space instead of only main road and lead to a shorter travel distance to the goal.
@@ IV EXPERIMENTS
In this section, we talk about the details of the implementation, comparisons, and ablation studies of this approach. The experiments are designed to demonstrate the benefits of our innovations:
@@
    Evaluate the novel End-to-end model, DTG, in mapless outdoor global navigation: We compare our approach, DTG, with state-of-the-art outdoor navigation algorithms, including MTG [liang2023mtg], NoMaD [sridhar2023nomad] and ViNT [shah2023vint] in both testing dataset and challenging outdoor real-world applications. The real-world experiment is achieved by combining the lower-level motion planner DWA [fox1997dynamic]. The details of the real-world experiment and the results are in the Appendix VI-A.
@@
    Evaluate the efficacies of Diffusion model, CRNN diffusion cells, and adaptive traversability loss for training: For each innovation, we conduct ablation studies to demonstrate the benefit and effect of the components. The diffusion mechanism is visualized in Appendix VI-B
@@ IV-A Implementation
As described in Section III, our major perceptual sensor is a 3D Velodyne Lidar (VLP-16), which has 16 channels and a 3Hz frequency. In this project, the inputs contains $C_{l}=3$ consecutive frames of Lidar and $C_{v}=20$ consecutive frames of historic velocities. The goal is set by converting the GPS value to meters. During training, the goals are randomly selected within 60 meters, and for testing, the goals are selected beyond 50 meters. The trajectory generator, DTG, generates $M=16$ waypoints in each trajectory. The voxelization radius of the PointCNN in the Perception Encoder is 0.08m in this approach. The training data is the same as MTG [liang2023mtg], which is collected by the Husky robot. Velocities are collected from an odometer with 10Hz frequency. To keep all the perceptive information in the same time period, we select consecutive 10 frames of velocities as input. During training, the ground truth trajectories are generated by the A* algorithm, with the shortest travel distance to the goal, and the ground truth trajectories are threshold by 15 meters from the robot’s position and each ground truth trajectory has 16 waypoints. The training and evaluation are processed in the computer with NVIDIA RTX A5000 GPU and an Intel Xeon(R) W-2255 CPU, and the real-world experiment is executed in the laptop with an Intel i7 CPU and one Nvidia GTX 1080 GPU. In the real-world experiment, the diffusion model generates trajectories in 5Hz and the traversability estimator estimates the traversability of the trajectories and keeps selecting the best trajectory during navigation. The network details of the architecture is in Appendix VI-C.
@@ IV-B Evaluation
In this section, we qualitatively and quantitatively evaluate DTG compared with different state-of-the-art methods and modified versions for ablation study. In the experiment, we compare DTG with ViNT [shah2023vint], NoMaD [sridhar2023nomad] and MTG [liang2023mtg]. Because the NoMaD and ViNT require a sequence of images from the start position to the goal, we run the robot in the environment first and collect the images. Each consecutive pair of images has a distance of around 1 meter. MTG generates multiple trajectories, and we use the same method as the experiment in MTG [liang2023mtg] to choose the best trajectory. The evaluation metrics include: Traversability, Distance Ratio, Inference Time, and Model Size.
@@
Distance Ratio: The distance ratio is to evaluate the future travel distance of the trajectory to the goal. The Distance Ratio measures the ratio of the travel distance of the trajectory w.r.t. the shortest travel distance from the robot’s current position to the goal. Therefore, given the trajectory length $|\hat{\bm{\tau}}|$, the travel distance $h_{c}$ from the robot’s position and the travel distance $h_{t}$ from the last waypoint $\mathbf{w}_{M}\in\hat{\bm{\tau}}$, the distance ratio is defined as Equation 12. The ratio is trajectory length independent, so we can use it to compare trajectory generators with different trajectory lengths.
@@
TABLE I: Quantitative Results: Our approach achieves $15\%$ improvement in distance ratio and around at least $7.6\%$ increase in traversability than ViNT and NoMaD. Our proposed CRNN achieves comparable traversability and distance ratio compared with U-Net, but our model is much faster and smaller.
@@
Comparisons: This section shows the efficacy of DTG for mapless outdoor global navigation. From Table I, we observe our approach, DTG, outperforms other SOTA approaches. Specifically, we achieve by at least $15\%$ in distance ratio and $12.6\%$ in traversability improvement compared with ViNT and NoMaD. NoMaD and ViNT only take RGB images, so the robot is not very robust in generating trajectories only in traversable areas. NoMaD is slightly better than ViNT w.r.t. traversability and inference time. Our approach outperforms MTG around $7\%$ in traversability and $16\%$ in distance ratio. The MTG has relatively good traversability, but cannot choose the trajectory with the best distance ratio, because in MTG the trajectories are generated by only comparing the straight distance to the goal instead of estimating the real travel distance of the trajectory. DTG also outperforms NoMaD and ViNT in inference time by 0.11s and 0.56s, but the model size is more than ViNT, NoMaD, and MTG.
@@
As shown in Figure 3 and 4, we provide the qualitative explanation of how our approach outperforms other SOTA methods. The top row shows the generated trajectory in the camera and the bottom row shows the trajectory in the traversability map, where Cyan color represents the generated trajectory and Yellow dots are the most ground truth trajectory with the shortest travel distance to the goal. From the two figures, we observe NoMaD and ViNT generate shorter trajectories than MTG or DTG. Figure 3 shows a challenging occluded environment around a corner, DTG can generate the trajectory aligned with the geometric shape of the traversable areas, but MTG and NoMaD do not perform well in challenging corner situations. In Figure 4, the goal is behind the building and the ground truth trajectory lies in a narrow passage. The scenario is challenging in terms of travel distance estimation. DTG can still generate a trajectory to the target in the narrow passage, while other approaches all choose trajectories in the wider main road.
@@
Ablation Study: To evaluate the capability of different components of our innovations, we compare DTG with the modified versions without traversability loss during training, changing our Conditional RNN to the regular U-Net model and changing the generative model from the diffusion model to CVAE [cvae]. From Table I, our DTG has the best heuristic compared with other ablation studies. The U-Net is much larger than our proposed CRNN with more than 989.51Mb and is also much slower than all other approaches, but because the U-Net model is larger than CRNN, it can encode information better and generates trajectories with slightly better traversability. Our novel model CRNN is faster and smaller than U-Net but we achieve comparable distance ratio and traversability in trajectory generation. The DTG/t shows the model without training traversability loss, and obviously, it has the worst traversability, whereas NoMaD has the same issue in training diffusion models. The CVAE model is smaller, but the generative capability is not as good as diffusion models. Its distance ratio and traversability of the generated trajectories are worse than DTG.
V Conclusion, Limitations and Future Work
@@
We present a novel end-to-end diffusion-based trajectory generator for mapless global navigation. We demonstrated the innovations of the end-to-end approach for mapless outdoor navigation tasks and also the efficacy of the different innovative components. We achieve at least $15\%$ improvement in distance ratio and $7\%$ improvement in traversability compared with other SOTA methods.
@@
There are also limitations of the approach, DTG. We do not estimate the confidence of the generated trajectories, because the trajectory generator generates trajectories in real-time, there should be some mechanism to smartly choose the best trajectory during navigation. For example, estimating the feasibility and confidence of the trajectory.

### 25 Mar 2024 v2

@@ III-C Training Strategy
According to the problem defined in Section III-A, we have two targets to achieve: 1. Reduce the distance between the generated trajectories and the ground truths, and 2. Minimize the portion of generated trajectories in non-traversable areas. Since our approach is an end-to-end model, we jointly train both targets with an adaptive training strategy:
@@
where $t\in\left\{1,...,N\right\}$ and $\epsilon_{d}\in\mathcal{N}(0,\mathbf{I})$. $\bm{\tau}_{gt}$ is the ground truth trajectory because optimizing VDM [vdm] boils down to predicting the original ground truth. SNR$(t)=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}$. $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ is defined the same as in [ho2020denoising]. To simplify the loss function, because SNR(t) is monotonically decreasing, $(\text{SNR}(t-1)-\text{SNR}(t))$ is always positive. We find similar training results after removing this term during training. Thus, the loss function is simplified as:
@@ III-C2 Adaptive Training of Diffusion Models
where $d(\cdot)$ calculates the distance between the current waypoint $w_{m}$ in $\hat{\bm{\tau}}_{t}$ and its nearest non-traversable area The distance function clips the value to $[0-1]$ meters. We sample t from the diffusion step buffer $\mathbf{b}_{t}$, which stores the currently available diffusion steps. As shown in Algorithm 1, $i$ is the index of training epochs and $N_{t}=10$ is the maximum step number to apply to the traversability loss. $l_{d}(t-1)$ calculate the average of $5$ last recorded diffusion loss values at step $t-1$. If the average loss value is smaller than the threshold $H_{d}$, we add this step to the step buffer $\mathbf{b}_{t}$. We increase the range of diffusion steps incrementally to adapt the training of traversability to the training of ground truth trajectories. Finally, we have the total loss function for DTG: $\mathcal{L}=\mathcal{L}_{d}+\beta\mathcal{L}_{t}$, where $\beta$ is a hyperparameter.
@@
Figure 3: Traversability Analysis in Challenging Occluded Environment: The top row shows the generated trajectories (red) in the camera view. The bottom row shows the top-down view of the traversability map. The cyan color represents the generated trajectories, and the yellow color represents the most heuristic trajectory to the goal. DTG can generate trajectories w.r.t. the geometric shape of the traversable areas, but other approaches cannot generate fully traversable trajectories; the non-traversable parts are marked by red circles.
@@
Figure 4: Travel-Distance Analysis in Challenging Narrow Passage Environment: The goal is behind the building. DTG generates a trajectory in a narrower space instead of the wide main road, leading to a shorter travel distance to the goal.
@@ IV EXPERIMENTS
In this section, we discuss the details of the implementation, comparisons, and ablation studies of this approach. The experiments are designed to demonstrate the benefits of our innovations:
@@
    Evaluate the novel end-to-end model, DTG, in maples outdoor global navigation: We compare our approach, DTG, with SOTA outdoor navigation algorithms, including MTG [liang2023mtg], NoMaD [sridhar2023nomad], and ViNT [shah2023vint] in both a testing dataset and challenging outdoor real-world scenarios. The real-world experiment is achieved by combining the lower-level motion planner DWA [fox1997dynamic]. The details of the real-world experiment and the results are in Appendix VI-A[supplements].
@@
    Evaluate the efficacies of Diffusion model, CRNN diffusion cells, and adaptive traversability loss for training: For each innovation, we conduct ablation studies to demonstrate the benefit and effect of the components. The diffusion mechanism is visualized in Appendix VI-B[supplements]
@@ IV-A Implementation
As described in Section III, our major perceptual sensor is a 3Hz 3D Velodyne Lidar (VLP-16) with 16 channels. The inputs of this approach contain $C_{l}=3$ consecutive frames of Lidar and $C_{v}=20$ consecutive frames of historic velocities. The goal is set by converting the GPS value to meters. During training, the goals are randomly selected within 60 meters; for testing, the goals are selected beyond 50 meters. The trajectory generator, DTG, generates $M=16$ waypoints in each trajectory. The voxelization radius of the PointCNN in the Perception Encoder is 0.08m in this approach. The training data is the same as MTG [liang2023mtg], which is collected by a Husky robot. Velocities are collected from a 10Hz odometer. To keep all the perceptive information in the same time period, we select consecutive 10 frames of velocities as input. During training, the ground truth trajectories, generated by the A* algorithm with the shortest travel distance to the goal, are thresholded 15 meters from the robot’s position, and have 16 waypoints each. The training and evaluation are processed in a computer with an NVIDIA RTX A5000 GPU and an Intel Xeon(R) W-2255 CPU, and the real-world experiment is executed on a laptop with an Intel i7 CPU and one Nvidia GTX 1080 GPU. In the real-world experiment, the diffusion model generates trajectories in 5Hz. The network details of the architecture are in Appendix VI-C[supplements].
@@ IV-B Evaluation
In this section, we qualitatively and quantitatively evaluate DTG compared with different state-of-the-art methods and modified versions for ablation study. In the experiment, we compare DTG with ViNT [shah2023vint], NoMaD [sridhar2023nomad], and MTG [liang2023mtg]. Because NoMaD and ViNT require a sequence of images from the start position to the goal, we run the robot in the environment first and collect the images. Each consecutive pair of images has a distance of around 1 meter. MTG generates multiple trajectories, and we use the same method as the experiment in MTG [liang2023mtg] to choose the best trajectory. The evaluation metrics include: Traversability, Distance Ratio, Inference Time, and Model Size.
@@
Distance Ratio: The distance ratio is to evaluate the future travel distance of the trajectory to the goal and it measures the ratio of the travel distance of the trajectory w.r.t. the shortest travel distance from the robot’s position to the goal. Therefore, given the trajectory length $|\hat{\bm{\tau}}|$, the travel distance $h_{c}$ from the robot’s position, and the travel distance $h_{t}$ from the last waypoint $\mathbf{w}_{M}\in\hat{\bm{\tau}}$, the distance ratio is defined as Equation 12. The ratio is trajectory length independent, so we can use it to compare trajectory generators with different trajectory lengths.
@@
TABLE I: Quantitative Results: Our approach achieves at least a $15\%$ improvement in distance ratio and around $7\%$ increase in traversability over other approaches. Our innovative components, CRNN, adaptive traversability training, and end-to-end diffusion-based generator also show effective improvement in global navigation task.
@@
Comparisons: This section shows the efficacy of DTG for mapless outdoor global navigation. From Table I, we observe our approach, DTG, outperforms other SOTA approaches. Specifically, we achieve by at least $15\%$ in distance ratio and $12.6\%$ in traversability improvement compared with ViNT and NoMaD. NoMaD and ViNT only take RGB images, so the robot is not very robust in generating trajectories only in traversable areas. NoMaD is slightly better than ViNT w.r.t. traversability and inference time. Our approach outperforms MTG around $7\%$ in traversability and $16\%$ in distance ratio. The MTG has relatively good traversability but cannot choose the trajectory with the best distance ratio because in MTG the trajectories are generated by only comparing the straight distance to the goal instead of estimating the real travel distance of the trajectory. DTG also outperforms NoMaD and ViNT in inference time by 0.11s and 0.56s, but the model size is bigger than ViNT, NoMaD, and MTG.
@@
As shown in Figures 3 and 4, we provide the qualitative explanation of how our approach outperforms other SOTA methods. The top row shows the generated trajectory in the camera and the bottom row shows the trajectory in the traversability map, where cyan represents the generated trajectory and yellow dots are the ground truth trajectories with the shortest travel distance to the goal. From the two figures, we observe NoMaD and ViNT generate shorter trajectories than MTG or DTG. Figure 3 shows a challenging occluded environment around a corner; DTG can generate the trajectory aligned with the geometric shape of the traversable areas, but MTG and NoMaD do not perform well in challenging corner situations. In Figure 4, the goal is behind the building and the ground truth trajectory lies in a narrow passage. The scenario is challenging in terms of travel distance estimation. DTG can still generate a trajectory to the target in the narrow passage, while other approaches all choose trajectories in the wider main road.
@@
Ablation Study: To evaluate the capability of different components of our innovations, we compare DTG with the modified versions without traversability loss during training, changing our Conditional RNN to the regular U-Net model and changing the generative model from the diffusion model to CVAE [cvae]. From Table I, our DTG has the best heuristic compared with other ablation studies. The U-Net is much larger than our proposed CRNN with more than 989.51Mb and is also much slower than all other approaches, but because the U-Net model is larger than CRNN, it can encode information better and generates trajectories with slightly better traversability. Our novel model CRNN is faster and smaller than U-Net, but we achieve comparable distance ratios and traversability in trajectory generation. The DTG/t shows the model without training traversability loss. Obviously, it has the worst traversability. The CVAE model is smaller, but the generative capability is not as good as diffusion models. Its distance ratio and traversability of the generated trajectories are worse than DTG.
V Conclusion, Limitations, and Future Work
@@
We present a novel end-to-end diffusion-based trajectory generator for mapless global navigation and demonstrate the innovations of the end-to-end approach and the efficacy of the different innovative components in both evaluation dataset and the real-world experiments. We achieve at least a $15\%$ improvement in distance ratio and a $7\%$ improvement in traversability over other SOTA approaches.
@@
There are also limitations of the approach, DTG. Because the trajectory generator generates trajectories in real-time, there should be some mechanism to smartly choose the best trajectory during navigation, e.g., estimating the feasibility and confidence of the trajectory.